# “You Might Like It”: How People Respond to Small Talk in Human-Robot Collaboration

### v1

@@ 4.4.2. Blame attribution
Overall, 10 participants blamed no one for the system announcement, 4 blamed the robot, 3 the team, and 3 themselves. Among those blaming the robot, only one cited its small talk, while the rest attributed the announcements to its slow speed. One participant who blamed the team stated, “…Maybe in terms of like the team chemistry, that was probably where it was lacking”(P16). Another explained how they thought the novelty of the conversation impacted the team’s performance, “I also get a little more distracted, because …this is just …a new thing …it’s something that I guess you would get used to if you’re working with it every day”(P10).
Those who blamed themselves acknowledged possible mistakes. Meanwhile, 15 participants discussed they were confused at what prompted the system announcements because they thought the team was performing efficiently. P1 explained how the announcement “felt like out of my control. So I was like, I don’t think there was anything else that could have been done.”
@@ 5. Discussion
We explored the incorporation of conversational elements such as small talk into industrial robots, especially those not inherently built for social interactions.
@@
While opinions on the robot’s conversational abilities vary, participants showed a willingness to engage in off-topic conversations initiated by the robot. In this section, we discuss our findings and offer design guidelines for robot small talk.
5.1. Robot-Initiated Small Talk: First Impression
Participants’ immediate impressions and engagements in small talk with the robot validate considering such dialogue for non-social robots. Our findings (section 4.3) suggest the majority of participants ($N=17$) positively received the robot’s conversation attempts, attesting to the potential of robots to engage meaningfully in small talk right from the outset. Yet, not every response was unequivocally positive. Some participants abstained from engagement ($N=2$) or posed a rhetorical question ($N=1$, see Figure 4 bottom). Such user attitudes, potentially anchored in pre-established expectations, might indicate their expectation for manipulator robots to adhere strictly to task-related interactions (dennler2023design, 12).
Participants’ task engagement revealed insights into their receptivity to unexpected robot behavior (small talk) for a otherwise non-social robot (section 4.3). While most participants ($N=13$) briefly glanced at the robot, suggesting a brief attention shift without disruption, the ”freeze” reactions of some ($N=4$) signaled a task disruption. This echos prior work noting that extended mutual eye-gaze in human-human exchanges often signals ”problematic interruptions” (degutyte2021role, 11). The robot’s initial conversation also evoked affective reactions. The smiles from many participants ($N=8$) suggest robots can spark positive emotions early on. Although a smile might also indicate pleasant ”surprise”, overt signs of surprise from others ($N=6$) emphasize the unexpectedness of small talk from a primarily task-oriented robot. Such shift in user attention and emotional state underscores the unexpectedness of the first question, likely due to the robot’s unforeseen conversational approach and a perceived mismatch with its physical embodiment.
In summary, the notable engagement and minimal task disruption despite surprise indicate that even if participants are surprised by the robot’s first question, it does not hinder their task. This emphasizes participants’ readiness to converse with robots, even during ongoing tasks and when the unexpected conversation is most pronounced. As participants adapt to the robot’s conversational style, diversions due to small talk might reduce further. Future research should investigate optimal small talk design to balance task efficiency and user enjoyment. Calibrating conversation frequency is vital to avoid minimize excessive talkativeness that distracts, or causes awkwardness due to prolonged silence.
5.2. Potential of Small Talk for Otherwise Non-Social Robots
5.2.1. User engagement: Small talk in purely functional tasks
While engaging in small talk with a robot arm manipulator during collaborative tasks, we observed a noticeable willingness to not only respond to the robot’s direct questions but also its comments that do not warrant a user response (section 4.1). For instance, in Figure 1, the participant comments (“You might like it.”) on the robot’s response (“No, I have not”) to her question (“if you’ve ever played that?”). While most of the responses by the participants were single sentenced, 12% of the responses were multiple sentences showing higher and meaningful engagement (section 4.2.1). Such level of user engagement is evident. In Figure 1, the participant’s response ”(chuckle) I think it’s a pretty classic game …They even made a movie about it.” goes beyond merely answering the robot’s query; it offers supplementary information, perhaps in anticipation of the robot’s genuine interest in the game. It is as if the participant was attempting to pique the robot’s curiosity, subtly suggesting beforehand, “You might like it”.
Although the desire to ask robot questions in return varied among participants, 45% of them posed personal questions to the robot, inquiring about its preferences or capabilities (see section 4.2.2 and Figure 4). Additionally, participants perceived the small talk positively; P20 mentioned, “it was cool to have somebody to talk to”, while P16 characterized the conversation with the robot as, “more personal and …less mechanical”. Thus, our exploratory study indicates user engagement and responsiveness to robot-initiated off-topic conversations, suggesting potential for integrating small talk to enhance user engagement in human-robot collaborative tasks that involve physical manipulation.
Potential 1: Small talk may boost meaningful engagements in human-robot interactions, even with non-anthropomorphic robots in purely functional scenarios.
5.2.2. Ice-breaking through small talk: building rapport and unveiling robot capabilities
Participants demonstrated curiosity about the robot’s functions, often posing small talk questions to probe its capabilities and background. For instance, as illustrated in Figure 4 top, one participant inquired, ”Hey robot, tell me more about yourself. I’m very interested. Like, you said you’re from Germany.” Another, referencing Figure 4 bottom, sought clarity on the robot’s primary function by asking, ”Aren’t you here just to help me with these pipes?” Such questions highlight the potential of employing conversational frameworks, even for non-social robots, to build common ground and mutual understanding in human-robot teams. By structuring initial interactions around themes such as ”getting to know your teammate”, robots can ease introductions, reducing potential barriers to seamless collaboration. Participant P7’s sentiment, “I’m curious to learn more about it” encapsulates this curiosity. Thus, small talk can serve as an initial icebreaker, fostering team unity and rapport. Over time, or in subsequent sessions, these conversations can become instrumental in communicating the robot’s capabilities, ensuring that users develops a comprehensive understanding of its functions. Furthermore, by incorporating task-related insights, tidbits, or safety tips into these dialogues, the robot can harness users’ social attention to support task performance.
Potential 2: Small talk can be an effective ice-breaker, fostering rapport and understandings of robot capabilities in collaborative settings.
5.3. Challenges and Design Guidelines for Robot Small Talk Framework
5.3.1. Dynamic and user-aware conversational framework
The conversational framework we employed for robot small talk demonstrated constraints. The design was rigid, causing interactions to feel scripted rather than organic as pointed out by P16: “the robots question feel kind of scripted or …inorganic so,…asking it questions, which is kind of like asking questions to a voice assistant.” The robot’s knowledge bank was confined to a limited scope of topics (by design). For instance, in the conversation with P15, after the robot’s question “By the way, do you like to travel?” , the participant responded in negation and requested next pipe: “No. Bring me a yellow pipe”. The next robot question (“What is the coolest place you have ever been to?”) was not well received by the participant. Despite their clear negation to the previous question indicating disinterest, the robot asked a follow-up question on the same topic, agitating (“Shut up!”) by the participant. Similarly, P18 responded to the robot question about traveling: “You know I don’t know and I don’t feel like talking about traveling right now.” Thus, rigid conversational frameworks fail to adapt to the evolving nature of human conversation, highlighting the future need for a more dynamic, expansive and user-aware conversational approach.
To increase conversational flexibility, small talk should reflect user interests, allowing them to initiate and engage in a broader range of topics. Leveraging Large Language Models (LLMs) can help tailor small talk for robots and expand a robot’s knowledge. However, they must be adapted to align with a robot’s physical limitations, especially when addressing questions about their origins and experiences such as, ”How were you created?”, or ”Have you traveled?” Thus, LLMs can enhance the conversational capabilities but should not be utilized without necessary modifications via prompt engineering or fine tuning to avoid confusion about a robot’s physical capabilities and realistic experiences.
Design Guideline 1: Robots should use dynamic, user-aware conversational frameworks supporting flexible small talk interactions beyond rigid topics or formats.
5.3.2. Balancing small talk with task performance
Engaging in small talk with people during physical tasks poses challenges, such as with the frequency of questions posed. For instance, initiating small talk during every user command or robot action might be excessive as expressed by a participant (P15) through a rhetorical question to the robot: “Why do you talk so much?” Over-engagement in non-task related conversations can be distracting for users (veletsianos2012learners, 51, 31), potentially jeopardizing the overall performance or efficiency of the collaboration. Therefore, it is imperative to strike a balance between fostering rapport and ensuring task performance. Small talk frequency and complexity should be adapted based on task intricacy and observed team performance. During critical tasks, or when user confusion or distraction arises, it might be wise to limit or halt small talk. Using error and confusion detection frameworks (as proposed in (stiber2023using, 47, 27)) can help robots gauge progress and performance. Using these frameworks, robots can determine the appropriate time, or whether at all, to engage in small talk.
Design Guideline 2: Robots should ensure small talk is context-sensitive, enhancing team rapport without impeding task completion.
5.3.3. Mental model mismatch: managing user expectations
Several participants were unaware of interactive possibilities with the robot as evident in section 4.2.2. Participant P12 comments and inquires, “That’s strange coming from a robot. Aren’t you here to just help me with these pipes?” (see Figure 4 bottom). This points towards a potential misalignment in their mental model of the robot’s capabilities. P3’s statement, ”From beginning to end, I always felt like it’s a robot [non-human form]”, further illustrates this misalignment. It raises a pivotal question: What conversational depth is appropriate for a robot manipulator that lacks human-like form and operates within constrained physical and cognitive boundaries?
Conversely, when robots participate in conversations, it might inflate people’s expectations beyond the robot’s actual capabilities. For instance, few participants expected the robot to be more capable, only to find the robot lacking, as P7 pointed out: “I mean, because …right now, artificial intelligence is so popular. I thought a robot [is] gonna be more intelligent.”
Such misalignment may be exacerbated by participants’ familiarity with advanced conversational AI tools, such as Alexa and ChatGPT, leading them to expect that the “social” robot possesses enhanced conversational abilities. Therefore, when a physical robot starts chatting, it may lead to misalignment in user expectations and robot capabilities, as P10 expressed: “Because once [robot] starts asking me stuff, then I was …oh, maybe it’s …kind of like GPT-esque, where …has further knowledge rather than just questions, but …then my expectation went back a little bit.” If the robot’s dialogue suggests a high degree of intelligence, any shortfall in meeting those expectations could lead to increased user disappointment. Even slight mismatches between the voice and appearance of robots have been shown to introduce artifacts in human-robot interactions (mcginn2019can, 29). Establishing common ground and rapport crucially relies on instilling an accurate mental model of the robot in individuals’ minds (kiesler2005fostering, 19). Thus, it is essential to consider how user expectations can be managed adequately while designing social characteristics for robots. The robot’s social cues may shape individuals’ perceptions; if misaligned initially, the robot’s early behaviors can correct these views and recalibrate common ground (kiesler2005fostering, 19).
Design Guideline 3: Intentionally calibrate user expectations by matching robot conversational cues with its capabilities, considering users’ familiarity with AI.

### v2

@@ 4.3.2. Blame Attribution
Overall, 10 participants blamed no one for receiving the negative system feedback, 4 blamed the robot, 3 the team, and 3 themselves. Of those blaming the robot, only one cited its small talk while the rest attributed the announcements to its slow movement speed. One participant who blamed the team stated, “…Maybe in terms of like the team chemistry, that was probably where it was lacking” (P16). Another explained how they thought the novelty of the conversation impacted the team’s performance, “I also get a little more distracted, because …this is just …a new thing …it’s something that I guess you would get used to if you’re working with it every day”(P10). Those who blamed themselves acknowledged possible mistakes. Meanwhile, 15 participants discussed they were confused about what prompted the system announcements because they thought the team was performing efficiently. P1 explained how the announcement “felt like out of my control. So I was like, I don’t think there was anything else that could have been done.”
4.3.3. Participant Responses to Robot Small Talk After Feedback
[…]
@@ 5. Discussion
We explored the incorporation of conversational elements such as small talk into industrial robots, especially those not inherently built for social interactions. While opinions on the robot’s conversational abilities vary, participants showed a willingness to engage in off-topic conversations initiated by the robot. In this section, we discuss our findings in relation to prior work, identify design considerations for future conversational interactions, and acknowledge the limitations of our current work alongside variables that warrant further exploration.
5.1. Social Conversations in Physical Human-Robot Collaboration
Prior research indicates that people in manufacturing settings value social interactions; specifically, robot operators engaged in small talk at times with fellow operators during their shifts (Cheon et al., 2022). As it is common for humans to partake in human-human interaction with their co-workers, it is an interesting observation that humans may find themselves talking to their robot partners. Additionally, a different set of robot operators revealed in interviews that they occasionally spoke to the robot and expressed a desire for the robot to exhibit more social behaviors (Sauppé and Mutlu, 2015).
When asking whether or not we should include conversational robot interactions in manufacturing settings, we have to recognize that people are inclined to speak with one another and with their robot counterparts. This tendency to speak to the robot aligns with the evidence discovered in this work (see Sec. 4.2.2). Most of our participants engaged in casual conversations with their robot assistant; they responded to robot small talk arising from a traditionally non-social robot during a non-social collaborative assembly task.
In our study, the participants who exhibited non-task-oriented conversations while interacting with the robot shared different incentives for engaging with it. One individual explained their engagement may have been due to reciprocity, “It initiated questions to me. So I was like, Oh, I guess I can talk to it… I mean … if something’s talking to me, I feel like I typically talk back” (P10). However, many participants presented personal questions to the robot due to their curiosity about its functionalities and background (see section 4.2.3). For instance, as illustrated in Figure 6 top, the participant inquired, “Hey robot, tell me more about yourself. I’m very interested. Like, you said you’re from Germany.” Another participant, referencing Figure 6 bottom, sought clarity on the robot’s primary function by asking, “Aren’t you here just to help me with these pipes?”. It seems as if the robot’s display of its verbal capabilities was an invitation for the user to communicate in return with a verbal response.
The participants’ willingness to advance the conversation beyond simple or short responses is demonstrated by their medium and multi-sentence responses, resulting in about 44% and 15% of their social responses that were not a question (see section 4.2.1). Users displayed a noticeable disposition to respond to the robot’s direct questions, provide supplemental information, and reply to the robot’s comments that did not warrant a user response. This level of user engagement is evident in Figure 1, where a participant’s response goes beyond merely answering the robot’s query, not only describing aspects of the discussed game (Tetris) but anticipating the robot’s genuine interest in the game by commenting (“You might like it.”) on the robot’s response (“No, I have not”) to her question (“if you’ve ever played that?”). When describing the game’s premise in additional detail, the participant not only made eye contact with the robots’ body but also used hand and arm gestures to supplement their explanation. We found this participant’s use of gestures a bit surprising, as the robot does not have a face, we did not place cameras on the robot itself, and we did not inform participants with details about the robot’s visual processing capabilities.
Overall, it appears that participants wanted to talk to the robot and provide more meaningful responses during their conversations. We thus are making a case for more social communication with robots during physical, collaborative settings; given the findings in the literature and our observations of meaningful interactions during this study, we argue for further exploration and for designing social conversations between humans and robots. Next, we will discuss our lessons learned and design considerations when creating future conversational interactions.
Table 5. Summary of variables that we believe may influence robot small talk in physical, collaborative contexts.
Theme Variables
User Experiences and Preconceived Notions Affecting Small Talk Preferences Acceptance of Small Talk (General)
[…]
Small Talk Characteristics Dynamic and Flexible
[…]
Task Context and Performance Balanced with Task Cognitive Demand
[…]
5.2. Lessons Learned and Key Takeaways
While our findings revealed that most individuals did engage with the robot in small talk, during post-study interviews, we learned that users had separate preferences regarding small talk in general and divergent attitudes toward the robot’s social behavior. Considering participants’ individual differences, we will discuss the various variables discovered that may impact how researchers design future small-talk interactions for robots that work with users in physical collaborative scenarios and other notable takeaways that emerged during our analysis. The main themes of these variables and takeaways are summarized in Table 5: user experiences and preconceived notions affecting small talk preferences, small talk characteristics, and task context and performance.
5.2.1. User Perceptions of Robots and their Effect on Small Talk Preferences
We discovered that users had different attitudes towards robots in general and robots that exhibited small talk during a collaborative task. While we asked participants about their prior experiences with technology and robots before the study, we did not ask for further details to better understand these past interactions. Thus, we were unaware of their preferences towards engaging in small talk with technology or any prior use of smart speakers for daily assistance. Our findings (section 4.1) suggested that most participants ($n=17$) responded positively to the robot’s conversation attempts; the smiles from many participants ($n=8$) after the robot’s first small talk question demonstrated positive emotions early on. One participant enjoyed conversing with their robot partner, explaining, “He asked me some questions about my, like leisure time, so I think he’s pretty delightful to work with… I’m happy to work with him” (P7). Another participant characterized the conversation with the robot as “more personal and …less mechanical” (P16). On the contrary, we found that two participants abstained from verbal engagement altogether, and one participant posed a rhetorical question to the robot immediately after the first small talk attempt (see Figure 6 bottom). Such user attitudes, potentially anchored in pre-established expectations, might indicate their expectation for manipulator robots to adhere strictly to task-related interactions (Dennler et al., 2023). As one of the participants who did not engage with the robot at all later explained, “It’s not a human, you don’t have a conversation” (P17). This user did not accept the idea of possibly conversing with the robot, as they refused to socially respond despite the numerous small talk initiation attempts made by the robot; they held firm in their conception of the robot as a non-social entity. The participant who responded to the robot’s first question with a rhetorical question communicated their surprise at the sociability of the robot, commenting and inquiring, “That’s strange coming from a robot. Aren’t you here to just help me with these pipes?” (see Figure 6 bottom). Additionally, this user’s tone throughout their small talk responses appeared to be filled with annoyance and sarcasm. This reaction points towards a potential misalignment in the user’s mental model of the robot’s capabilities. Yet, while this participant may not have initially associated robots with social capabilities, they still interacted with the robot, requesting further information about its role in the task.
However, a different participant stopped engaging in small talk with the robot altogether after being dissatisfied with its conversational capabilities. In the conversation with P15, after the robot’s question “By the way, do you like to travel?”, the participant responded in negation and requested the next pipe: “No. Bring me a yellow pipe”. The next robot question (“What is the coolest place you have ever been to?”) was not well received by the participant. Despite their clear negation to the previous question indicating disinterest, the robot asked a follow-up question on the same topic, agitating (“Shut up!”) the participant. When asked about the robot’s conversational capabilities, the participant explained they concluded that “the robot is not going to talk like a human being” (P15). Thus, the robot’s ability to adapt the conversation to topics of interest to the user may be crucial to maintaining their engagement. Other participants expressed how they felt that the robot’s questions were “random” (P26), the robot’s responses were “super generic” (P20), and one hated how the robot tried to talk to them about “unrelated topics” (P37). Having more knowledge of the users’ general interests, past experiences with technology, and current attitudes towards the robot’s capability to engage in small talk could help researchers better comprehend how individuals will accept a physically collaborative robot engaging in small talk ( Table 5). It could be helpful to devise metrics to not only determine which users will likely want to work with robots but also who may want to engage and participate in small talk. Factors such as user’s personality (Esterwood and Robert, 2022; Esterwood et al., 2021), cultural background, and country of origin (Mak and Chui, 2013; Rings, 1994; Yang, 2012) may play an important role in understanding user attitudes and preferences for small talk (Table 5).
5.2.2. Considerations for Small Talk Characteristics and Attribute
Engaging in small talk with people during physical tasks poses challenges, such as with the frequency of questions posed. For this study, the robot initiated conversations based on topics commonly used in small talk (Holmes, 2005), but it was programmed to occur during every user command or robot action, which might be perceived as excessive. One participant exclaimed through a rhetorical question to the robot: “Why do you talk so much?” (P15). Another participant expressed how the conversation felt too much for them as well, “Because it was like, more than I talk to my colleagues at work when I’m trying to get a task done” (P13). Over-engagement in non-task-related conversations can be distracting (Veletsianos, 2012; Methot et al., 2021), potentially lowering users’ focus on completing collaborative tasks. Additionally, it appeared that some individuals’ perceptions of the robot were influenced by the ongoing conversation; the rigidness of the robot conversation’s design led interactions to feel unnatural: “the robot’s question feel kind of scripted or …inorganic so,…asking it questions, which is kind of like asking questions to a voice assistant”. They further elaborated: “the questions were probably like pre-programmed or something. So, it doesn’t feel like having small talk with an actual human teammate” (P16). Despite one participant stating how “the conversation part makes it more like human-like”, they still acknowledged how the robot’s conversational abilities needed some changes. They further explained, “If the conversation part was, like, improved a little more, it’d be cool to have that” (P10). These participants raised valid concerns as there is a need to make the system more fluid and less pre-programmed, as the conversational structure we employed for robot small talk demonstrated constraints; the robot’s knowledge bank was confined to a limited scope of topics (by design). Just as users’ personalities may differ, so could their interests; the inability to cater small talk topics to the user could negatively impact their interaction experience. Thus, rigid conversational structures fail to adapt to the evolving nature of human conversations, highlighting the future need for a more dynamic, expansive, and user-aware conversational approach (Table 5).
To increase conversational flexibility, robot small talk should reflect user interests, allowing them to initiate and engage in a broader range of topics. Leveraging large language models (LLMs) could allow for more adaptable and flexible robot conversations by tailoring small talk to each individual user, thus expanding a robot’s knowledge. Future work could investigate how LLMs could enable robots with the ability to regulate small talk frequency, timing, duration, and complexity to promote task focus as needed (Table 5).
However, LLMs must be adapted via prompt engineering or fine-tuning to align with a robot’s physical limitations, especially when addressing questions about their origins and experiences, such as, “How were you created?” or “Have you traveled?” During the post-study interviews, one participant admitted, “I mean, because …right now, artificial intelligence is so popular. I thought a robot [is] gonna be more intelligent.” They further elaborated, “I just like originally wasn’t expecting it to be talking to me like that. I thought it was more just like a commands thing. So I was like, a little surprised. And I realized it had like, a little more to it than I thought. And then I think I like overshot with my expectation” (P7). We speculate that this perspective may have emerged due to the individual being familiar with technology and have a background in STEM, thus researchers need to consider how to avoid inflating people’s expectations of a robot’s capabilities in future interactions. While LLMs may enhance future robots’ conversational capabilities, they should not be utilized without necessary adjustments to avoid confusion about a robot’s physical capabilities and realistic experiences. A robot’s conversational capabilities need to be studied further, considering specifically that we may not want to increase user expectations of robots’ capabilities. At the same time, personal factors such as familiarity with newer technology, such as LLMs, may cause people to expect conversations to be more dynamic.
5.2.3. Effect of Task Context on Small Task Design
In our experimental setup, we chose a repetitive task that was physical and not very cognitively demanding as a starting point to simulate day-to-day simple assembly tasks. Participants viewed the utter nature of the task as ultimately boring: “very mundane” and “simple” (P20). Another explained: “I don’t know if you can, …make building these …super enjoyable, I guess or …fun” (P10). However, the robot small talk was able to leave a positive impact, with one participant stating how “the conversation did definitely …make the process more enjoyable” (P6). Another person explained how they developed excitement about the task: “Since the robot, …started the conversation, I think it [the task] was kind of, …more exciting and fun” (P9).
Even with simple experimental tasks, participants were slightly distracted after the introduction of robot small talk. One participant explained: “I felt distracted when it was asking me questions” (P12). Tasks that are more cognitively demanding, such as those requiring slight memorization or additional reading and visual processing, might be further impacted if small talk is introduced and causes slight user distractions. Moreover, many tasks may not have a static cognitive load (Raney, 1993; Gwizdka, 2010). Therefore, small talk should be further adjusted based on the cognitive demand of tasks and team performance; it should be limited during critical tasks or user confusion. It is still unclear to what extent small talk impacted team performance quantitatively. Therefore, future work should include additional quantification metrics to measure the distractions caused by small talk in tasks with varying cognitive loads (Table 5).
While some users expressed being slightly distracted by the robot’s small talk, the novelty of the interactive system could have impacted the users’ overall experience, as our robot’s small talk was initially new to all users. Therefore, a learning effect could have been possible, as participants were only beginning to learn how to balance small talk with the task. This was further highlighted when P10 explained: ”this is just like a new thing. And it’s a little bit distracting at first … I guess you would get used to it if you’re working with it every day or something”. As we observed, participants adjusted their strategy from responding to robot small talk before requesting the pipe, to requesting the pipe before responding to small talk after negative feedback. This shows a learning curve upon the realization that the team performance is not optimum; additional training and familiarity with the system could minimize the effects of a learning curve (Anzanello and Fogliatto, 2011).
Additionally, the novelty of such a system could have prompted more user interaction (Reimann et al., 2023). As previously discussed, participants were motivated to initiate questions due to their curiosity about the robot’s preferences and behaviors (section 5.1). However, once participants become used to such a system, their engagement frequency could decrease as their curiosity is satisfied and the novelty settles over time (Serholt and Barendregt, 2016; Leite et al., 2013).
In this experimental setup, the robot’s slow speed (a design choice for user safety (Michalos et al., 2015)) may have impacted users’ notions of blame after the feedback. Nevertheless, only one participant blamed the robot’s small talk for the teams’ critiqued performance (see section 4.3.2). This shows that even in the presence of small talk, users do not immediately assign sole blame to the robot for the team’s “poor performance.”
Nevertheless, it is unknown how user distractions and compromised team performance due to small talk would translate over time across longer or repeated interactions. While the rate of conversation extension might have decreased slightly right after the first negative feedback announcement, users continued to engage with the robot socially (see Sec. 4.2.3). Additional quantitative metrics are needed to truly determine if small talk impacted team performance, the quality of the collaboration, and how user behavior around small talk changes over time, especially as workers are expected to engage in tasks for prolonged periods daily (Table 5). Moreover, more exploration is needed to understand how to alter human-robot small talk dynamics to balance engagement and optimize team performance, especially for tasks that require varying levels of cognitive effort.